# Improved upper bounds on key invariants of Erdős-Rényi numerical semigroups

### v2

@@ 1. Introduction
For these and many other definitions and properties relating to numerical semigroups, we refer the reader to Rosales and García-Sánchez’s book [RGS+09]. Numerical semigroups arise in toric geometry since their semigroup algebras are exactly the coordinate rings of affine monomial curves. For a range of applications of numerical semigroups, see Assi, D’Anna, and García-Sánchez’s book [ADGS20].
@@
Let $p<24e^{-8}$. Then,
$\lim_{M\to\infty}\mathrm{E}[\operatorname{e}(\mathcal{S}(M,p))]\leq 7000\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}\text{and}$
$\lim_{M\to\infty}\mathrm{E}[\operatorname{g}(\mathcal{S}(M,p))]\leq\lim_{M\to\infty}\mathrm{E}[\operatorname{F}(\mathcal{S}(M,p))]\leq\frac{5000}{p}\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}$
In particular, for both the Frobenius number and the embedding dimension, we reduce the ratio between the lower and upper bounds from order $1/p$ to order $\left(\log(1/p)\right)^{3}$ at the cost of large constant factors and an assumption that $p$ is sufficiently small. The same results hold for the genus because it is trivially bounded by the Frobenius number.
@@
Our proof strategy is to separate the two processes as follows. First, we show that with high probability, both a prime $q$ of an appropriate size (depending on $p$) and a set $\mathcal{A}$ of roughly $\log(q)$ elements less than $q$ are selected as generators of $\mathcal{S}$. The set $\mathcal{A}$ can be interpreted as a random subset of the cyclic group $\mathbb{Z}_{q}$, and we prove a result (Theorem 2.1) about $k$-fold sums of such subsets that may be of independent interest in additive number theory. We use this result to show that with high probability, the Frobenius number is bounded by roughly $\log^{2}(q)$. We use simpler methods to bound the conditional expectation in the case that any of the high-probabilility events described above fail to occur. Finally, we adapt most of the same ideas to prove the bound on the expected embedding dimension.
The organization of the rest of the paper is as follows. In Section 2 we prove Theorem 2.1 via the second moment method. In Section 3 we apply Theorem 2.1 to show that in the model $\mathcal{S}(p)$, the Frobenius number is bounded by an appropriate function of $\frac{1}{p}$ with high probability. In Section 4 we complete the proof of Theorem 1.6. In Section 5 we present the results of experiments that suggest that the true expected values of the Frobenius number and embedding dimension lie asymptotically strictly between the lower bounds and our new upper bounds. More precisely, in both cases, the experiments point to exactly one factor of $\log(1/p)$ instead of zero (as in the lower bounds) or three (as in the upper bounds.) We end with some suggestions about how the upper bounds might be further improved.
@@ 2. Sumsets of Random Subsets of $\mathbb{Z}_{q}$
$\mathrm{Pr}\left(\left(b\log_{2}q\right)\mathcal{A}\neq\mathbb{Z}_{q}\right)\leq\frac{2b\log_{2}q+3}{q^{b-2}}.$
@@
    Our proof uses combinatorial bounds on binomial coefficients in order that the inequalities hold for all $q$. Asymptotically as $q$ tends to $\infty$ one could instead use Stirling’s formula to obtain the improved bound
    $\mathrm{Pr}\left(\left(b\log_{2}q\right)\mathcal{A}\neq\mathbb{Z}_{q}\right)\in O\left(\frac{\log q}{q^{2b-2}}\right).$
@@
3. Apéry sets modulo a selected prime generator
@@
Accordingly, the principal part of our strategy to prove Theorem 1.6 will be to use Theorem 2.1 to show that all of the elements of the Apéry set of some prime $q$ in $\mathcal{S}$ is bounded by a certain function of $\frac{1}{p}$ with high probability.
Let $b,c,d$ be positive constants to be determined later and set
$f(p)=\frac{d}{p}\left(\ln\left(\frac{d}{p}\right)\right)^{2}.$
Consider the following events for a semigroup $\mathcal{S}$ chosen via the random model $\mathcal{S}(p)$.
  • Define $D_{1}$ to be the event that at least one prime number between $f(p)+1$ and $cf(p)$ is selected as a generator of $\mathcal{S}$.
  • Assuming $D_{1}$, let $q$ be the largest prime selected in this range. Define $D_{2}$ to be the event that at least $2b\log_{2}q$ generators are selected between 1 and $q-1$.
  • Assuming $D_{1}$ and $D_{2}$, let $\mathcal{S}^{\prime}$ be the subsemigroup of $\mathcal{S}$ generated by a random subset of $2b\log_{2}q$ of the generators of $\mathcal{S}$ between 1 and $q$ along with $q$ itself if it is not already in the random subset. Let $D_{3}$ be the event that the largest element of the Ápery set of $\mathcal{S}^{\prime}$ with respect to $q$ is at most $bq\log_{2}q$.
If all of these events occur, then we can bound the Frobenius number by
(4) $\operatorname{F}(\mathcal{S})\leq\operatorname{F}(\mathcal{S}^{\prime})\leq bq\log_{2}q\leq bcf(p)\log_{2}(cf(p))=\frac{bcd}{p}\left(\ln\left(\frac{d}{p}\right)\right)^{2}\log_{2}\left(\frac{cd}{p}\left(\ln\left(\frac{d}{p}\right)\right)^{2}\right).$
3.1. Choosing an appropriate prime generator
To analyze the event $D_{1}$, we will apply a classic quantitative version of the Prime Number Theorem. Here $\pi(x)$ denotes the number of primes less than or equal to $x$.
Lemma 3.1.
Let $c\geq 7$ and $a\geq e^{c}$. Then $\pi(ca)-\pi(a)\geq\frac{a(c-7)}{\ln(ca)}.$
@@ 3.1. Choosing an appropriate prime generator
Rosser [Ros41] showed that for all $x\geq 55$,
@@
Here $ca\geq a\geq e^{7}>55$, so
$\pi(ca)-\pi(a)$ $\geq\frac{ca}{\ln(ca)+2}-\frac{a}{\ln a-4}$
$=a\left(\frac{c}{\ln c+\ln a+2}-\frac{1}{\ln a-4}\right)$
$=a\left(\frac{c\ln a-4c-\ln c-\ln a-2}{(\ln a)^{2}+\ln c\ln a-2\ln a-4\ln c-8}\right)$
$\geq a\left(\frac{c\ln a-4c-\ln c-\ln a-2}{(\ln a)^{2}+\ln c\ln a}\right)$
$=\frac{a}{\ln a+\ln c}\left(c-\frac{4c}{\ln a}-\frac{\ln c}{\ln a}-1-\frac{2}{\ln a}\right)$
$\geq\frac{a}{\ln a+\ln c}\left(c-4-1-1-2\right)$
$=\frac{a(c-7)}{\ln(ca)}.\qed$
Proposition 3.2.
Let $c\geq 7$ and $d>0$. For any probability $p\leq de^{-c}$, we have
$\mathrm{Pr}[\lnot D_{1}]\leq\left(\frac{p}{d}\right)^{\frac{d(c-7)}{3}}.$
Proof.
Let $N=\pi(cf(p))-\pi(f(p))$. We have $f(p)\geq\frac{d}{p}\geq e^{c}$, so we can apply Lemma 3.1 to obtain
$N$ $\geq\frac{\frac{d}{p}\left(\ln\left(\frac{d}{p}\right)\right)^{2}(c-7)}{\ln\left(\frac{cd}{p}\left(\ln\left(\frac{d}{p}\right)^{2}\right)\right)}$
$=\frac{d(c-7)}{p}\left(\frac{\left(\ln\left(\frac{d}{p}\right)\right)^{2}}{\ln\left(\frac{d}{p}\right)+\ln\left(2\ln\left(\frac{d}{p}\right)\right)+\ln c}\right)$
$>\frac{d(c-7)}{p}\left(\frac{\left(\ln\left(\frac{d}{p}\right)\right)^{2}}{3\ln\left(\frac{d}{p}\right)}\right)$
$=\frac{d(c-7)\ln\left(\frac{d}{p}\right)}{3p}$
where the last inequality holds since $x\geq 2\ln x$ for all $x\geq 0$ and also $\ln c<c<\ln\left(\frac{d}{p}\right)$ by hypothesis. Thus the probability of no prime being selected in this range is
$(1-p)^{N}\leq e^{-pN}\leq e^{\frac{-d(c-7)\ln p}{3}}=\left(\frac{p}{d}\right)^{\frac{d(c-7)}{3}}.\qed$
3.2. Choosing enough small generators
We now assume that the event $D_{1}$ holds and define $q$ to be the largest prime between $f(p)$ and $cf(p)$ that is selected as a generator. Let $G$ be the number of selected generators between 1 and $q-1$.
@@ 3.2. Choosing enough small generators
The reason this may not be obvious is that $q$ itself is a random variable that depends on which generators were selected. Indeed, if we had defined $q$ to be the smallest prime or a random prime rather than the largest prime among the generators between $f(p)+1$ and $cf(p)$, then this statement would not hold.
@@
$\mathrm{E}[G_{q_{i}}|q=q_{j}]=\frac{\mathrm{Pr}\left[(\textup{select }q_{i})\land(q=q_{j})\right]}{\mathrm{Pr}[q=q_{j}]}=\frac{p^{2}(1-p)^{n-j}}{p(1-p)^{n-k}}=p$
@@
Let $d\geq 4b$, $c\geq 7$, and $p\leq de^{-c}$. Then $\mathrm{Pr}[\lnot D_{2}|D_{1}]\leq\left(\frac{p}{d}\right)^{\frac{d}{8}\ln\left(\frac{d}{p}\right)}$.
@@
$\mathrm{E}[G]=p(q-1)\geq pf(p)=d\left(\ln\left(\frac{d}{p}\right)\right)^{2}.$
Now the hypothesis that $p\leq de^{-c}$ is equivalent to $c\leq\ln\left(\frac{d}{p}\right).$ In particular, $\frac{d}{p}\geq e^{c}\geq e^{7}>20$. Using these observations along with the hypothesis that $d\geq 4b$, the choice of $q$, and the fact that $\ln\left(x(\ln x)^{3}\right)\leq\left(\ln x\right)^{2}$ for $x>20$, we calculate that
$2b\ln q$ $\leq 2b\ln\left(\frac{cd}{p}\left(\ln\left(\frac{d}{p}\right)\right)^{2}\right)$
$\leq 2b\ln\left(\frac{d}{p}\left(\ln\left(\frac{d}{p}\right)\right)^{3}\right)$ $\leq 2b\left(\ln\left(\frac{d}{p}\right)\right)^{2}$
$\leq\frac{d}{2}\left(\ln\left(\frac{d}{p}\right)\right)^{2}$ $\leq\frac{E[G]}{2}.$
We apply the Chernoff bound
@@
with $\lambda=\frac{\mathrm{E}[G]}{2}$. Thus
$\mathrm{Pr}[\lnot D_{2}|D_{1}]\leq\mathrm{Pr}\left[G\leq\frac{\mathrm{E}[G]}{2}\right]\leq e^{-\frac{\mathrm{E}[G]}{8}}\leq e^{\frac{-d}{8}\left(\ln\left(\frac{d}{p}\right)\right)^{2}}=\left(\frac{p}{d}\right)^{\frac{d}{8}\ln\left(\frac{d}{p}\right)}.\qed$
3.3. Completing the Apéry set
We now assume that events $D_{1}$ and $D_{2}$ both hold, so we can apply Theorem 2.1.
@@ 3.3. Completing the Apéry set
For any $b>0$, $Pr[\lnot D_{3}|D_{1}\land D_{2}]\leq\frac{2b\log_{2}\left(cf(p)\right)+3}{\left(f(p)\right)^{b-2}}$.
@@
Using Theorem 2.1 along with the fact that $f(p)\leq q\leq cf(p)$, we calculate that
$Pr[\lnot D_{3}|D_{1}\land D_{2}]$ $\leq\frac{2b\log_{2}q+3}{q^{b-2}}$
$\leq\frac{2b\log_{2}\left(cf(p)\right)+3}{\left(f(p)\right)^{b-2}}\qed$
@@ 4. Proof of the Main Theorem
To prove the expected value bounds in Theorem 1.6, we will need an alternative bound on the Frobenius number to cover the case that one or more of the events $D_{1},D_{2},D_{3}$ do not occur. It is well known ([RGS+09, Proposition 2.13]) that for any coprime numbers $n_{1},n_{2}$, the Frobenius number of the semigroup $\langle n_{1},n_{2}\rangle$ is exactly $(n_{1}-1)(n_{2}-1)-1$. In particular, as soon as two consecutive numbers $12n,2n+$ 1 (the first one even) are selected as generators of $\mathcal{S}$, then
@@
$\mathrm{E}[\operatorname{F}(\mathcal{S})|\operatorname{F}(\mathcal{S})\geq u]\leq\frac{8}{p^{4}}+\frac{4u}{p^{2}}+u^{2}.$
@@
Let $N$ be the smallest number such that $N>u/2$ and such that both $2N$ and $12N+$ 1 are selected as generators. Since $p$ is constant, this random variable is finite with probability one. Also, $N$ depends only on the selection of generators greater than $u$. On the other hand, the event that $\operatorname{F}(\mathcal{S})\geq u$ depends only on the generators less than or equal to $u$, so $N$ is independent of this event. Thus
$\mathrm{E}[\operatorname{F}(\mathcal{S})|\operatorname{F}(\mathcal{S})\geq u]$ $\leq\mathrm{E}[\operatorname{F}\left(\langle 2N,2N+1\rangle\right)]$
$\leq\mathrm{E}[4N^{2}]$
$=\sum_{n=\frac{u}{2}+1}^{\infty}4n^{2}\left(1-p^{2}\right)^{n-\frac{u}{2}}p^{2}$
@@
$=\frac{8}{p^{4}}+\frac{4u}{p^{2}}+u^{2}\qed$
Let $F$ and $Z$ be random variables that represent the Frobenius number and the embedding dimension, respectively, of a numerical semigroup in $\mathcal{S}(p)$. In order to simultaneously bound the expected values of both $F$ and $Z$, we choose the parameter values $b=6$, $c=8$, and $d=24$. Let
$u(p)=bcf(p)\log_{2}(cf(p))=48f(p)\log_{2}(8f(p))$
so that as in (4), $\operatorname{F}(\mathcal{S})\leq u(p)$ whenever the events $D_{1}$, $D_{2}$, and $D_{3}$ all hold.
Lemma 4.2.
For the parameter values $b=6$, $c=8$, $d=24$ we have
$\mathrm{Pr}[\lnot(D_{1}\land D_{2}\land D_{3})]\leq 3\left(\frac{p}{24}\right)^{4}.$
Proof.
By Proposition 3.2 we have
$\mathrm{Pr}[\lnot D_{1}]\leq\left(\frac{p}{24}\right)^{8}\leq\left(\frac{p}{24}\right)^{4}.$
Since $p<24e^{-8}$, we have $\ln\left(\frac{24}{p}\right)>8$ and so by Proposition 3.4,
$\mathrm{Pr}[\lnot D_{2}|D_{1}]\leq\left(\frac{p}{24}\right)^{3\ln\left(\frac{24}{p}\right)}<\left(\frac{p}{24}\right)^{24}<\left(\frac{p}{24}\right)^{4}.$
Finally, by Corollary 3.5,
$\mathrm{Pr}[\lnot D_{3}|D_{1}\land D_{2}]$ $\leq\frac{2\cdot 6\log_{2}\left(8\left(\frac{24}{p}\right)\left(\ln\left(\frac{24}{p}\right)\right)^{2}\right)+3}{\left(8\left(\frac{24}{p}\right)\left(\ln\left(\frac{24}{p}\right)\right)^{2}\right)^{4}}$
$\leq\frac{12}{8^{4}}\left(\frac{p}{24}\right)^{4}\left(\frac{\log_{2}\left(8\cdot\left(\frac{24}{p}\right)\left(\frac{24}{p}\right)^{2}\cdot 8\right)}{\left(\log_{2}\left(\frac{24}{p}\right)\right)^{8}}\right)$
$\leq\left(\frac{p}{24}\right)^{4}\frac{\log_{2}\left(\left(\frac{24}{p}\right)^{4}\right)}{\log_{2}\left(\left(\frac{24}{p}\right)^{8}\right)}$
$\leq\left(\frac{p}{24}\right)^{4}.$
The result follows by the union bound. ∎
@@
We have
$u(p)$ $=48f(p)\log_{2}(8f(p))$
$=\frac{1152}{p}\left(\ln\left(\frac{24}{p}\right)\right)^{2}\log_{2}\left(8\cdot\frac{24}{p}\left(\ln\left(\frac{24}{p}\right)\right)^{2}\right)$
$<\frac{1200}{p}\left(\log_{2}\left(\frac{24}{p}\right)\right)^{2}\log_{2}\left(\frac{24}{p}\cdot\frac{24}{p}\left(\frac{24}{p}\right)^{2}\right)$
$=\frac{4800}{p}\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}.$
Then by Lemma 4.1 and the fact that $\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}<\frac{15}{p}$ for $p<24e^{-8}$, we have
$\mathrm{E}[F|F\geq u(p)]$ $\leq\frac{8}{p^{4}}+\frac{4\cdot 4800}{p^{3}}\left(\log_{2}\left(\frac{24}{p}\right)\right)^{3}+\frac{4800^{2}}{p^{2}}\left(\log_{2}\left(\frac{24}{p}\right)\right)^{6}$
(6) $<\frac{8}{p^{4}}+\frac{4\cdot 4800\cdot 15}{p^{4}}+\frac{4800^{2}*15^{2}}{p^{4}}$
$<\frac{6\cdot 10^{9}}{p^{4}}.$
We can now complete our proof by conditioning on whether or not the Frobenius number is bounded by $u$. For the Frobenius number itself, we apply Lemma 4.2 along with (4), (6), and the fact that $p\leq 24e^{-8}<\frac{1}{100}$ to obtain

### v3

@@ Abstract.
Keywords: random numerical semigroup, Frobenius number, probabilistic method, sumset
@@ 1. Introduction
For these and many other definitions and properties relating to numerical semigroups, we refer the reader to Rosales and García-Sánchez’s book [RGS09]. Numerical semigroups arise in toric geometry since their semigroup algebras are exactly the coordinate rings of affine monomial curves. For a range of applications of numerical semigroups, see Assi, D’Anna, and García-Sánchez’s book [ADGS20].
@@
As $p$ tends to zero,
$\mathrm{E}[\operatorname{e}(\mathcal{S}(p))]=O\left(\left(\ln\left(\frac{1}{p}\right)\right)^{3}\right)\text{and}$
$\mathrm{E}[\operatorname{g}(\mathcal{S}(p))]\leq\mathrm{E}[\operatorname{F}(\mathcal{S}(p))]=O\left(\frac{1}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{3}\right).$
In particular, for both the Frobenius number and the embedding dimension, we reduce the asympototic ratio between the lower and upper bounds from order $1/p$ to order $\left(\log(1/p)\right)^{3}$. The same holds for the genus because it is trivially bounded by the Frobenius number.
@@
Our proof strategy is to separate the two processes as follows. First, we show that with high probability, both a prime $q$ of an appropriate size (depending on $p$) and a set $\mathcal{A}$ of roughly $\log(q)$ elements less than $q$ are selected as generators of $\mathcal{S}$. The set $\mathcal{A}$ can be interpreted as a random subset of the cyclic group $\mathbb{Z}_{q}$, and we prove a result (Theorem 2.1) about $k$-fold sums of such subsets that may be of independent interest in additive number theory. We use this result to show that with high probability, the Frobenius number is bounded by roughly $\log^{2}(q)$. We use simpler methods to bound the conditional expectation of the Frobenius number in the case that the high-probabilility events described above fail to occur. Finally, we adapt most of the same ideas to prove the bound on the expected embedding dimension.
The organization of the rest of the paper is as follows. In Section 2 we prove Theorem 2.1 via the second moment method. In Section 3 we apply Theorem 2.1 to show that in the model $\mathcal{S}(p)$, the Frobenius number is bounded by an appropriate function of $\frac{1}{p}$ with high probability. In Section 4 we complete the proof of Theorem 1.6. In Section 5 we present the results of experiments that suggest that the true asymptotics of the expected Frobenius number and expected embedding dimension lie strictly between the lower bounds and our new upper bounds. More precisely, in both cases, the experiments point to exactly one factor of $\log(1/p)$ instead of zero (as in the lower bounds) or three (as in the upper bounds.) We end with some suggestions about how the upper bounds might be further improved.
@@ 2. Sumsets of Random Subsets of $\mathbb{Z}_{q}$
$\mathrm{Pr}[\left(b\log_{2}q\right)\mathcal{A}\neq\mathbb{Z}_{q}]\leq\frac{2b\log_{2}q+3}{q^{b-2}}.$
@@
    Our proof uses combinatorial bounds on binomial coefficients in order that the inequalities hold for all $q$. As $q$ tends to $\infty$ one could instead use Stirling’s formula to obtain the improved asymptotic bound
    $\mathrm{Pr}\left(\lceil b\log_{2}q\rceil\mathcal{A}\neq\mathbb{Z}_{q}\right)\in O\left(\frac{\ln q}{q^{2b-2}}\right).$
@@
3. Apéry sets and a typical generation process
In this section we will describe a series of events that imply a bound on the expected Frobenius number and expected dimension of a numerical semigroup chosen via the model $\mathcal{S}(p)$, and then bound the probability that any of these events fail to occur. The key to this bound will be Theorem 2.1.
@@ 3. Apéry sets and a typical generation process
Define
$f(p)=\frac{1}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{2}$
and consider the following events for a semigroup $\mathcal{S}$ chosen via the random model $\mathcal{S}(p)$.
  • Define $D_{1}$ to be the event that at least one prime number between $f(p)+1$ and $6f(p)$ is selected as a generator of $\mathcal{S}$.
  • Assuming $D_{1}$, let $q$ be the largest prime selected in this range. Define $D_{2}$ to be the event that $D_{1}$ occurs and also that at least $12\log_{2}q$ generators are selected between 1 and $q-1$.
  • Assuming $D_{2}$, let $\mathcal{S}^{\prime}$ be the subsemigroup of $\mathcal{S}$ generated by a random subset of $12\log_{2}q$ of the generators of $\mathcal{S}$ between 1 and $q$, along with $q$ itself. Let $D_{3}$ be the event that $D_{2}$ occurs and also that the largest element of the Apéry set of $\mathcal{S}^{\prime}$ with respect to $q$ is at most $6q\log_{2}q$.
If $D_{3}$ holds, then the semigroup $\mathcal{S}^{\prime}$ defined in $D_{3}$ contains $\mathcal{S}$. Using (3) and the upper bound on $q$, we can bound the Frobenius number of $\mathcal{S}$ by
(4) $\operatorname{F}(\mathcal{S})\leq\operatorname{F}(\mathcal{S}^{\prime})\leq 6q\log_{2}q\leq 36f(p)\log_{2}(6f(p))\leq K\left(\frac{1}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{3}\right).$
for a sufficiently large constant $K$.
To bound the probability that $D_{3}$ does not hold, we begin by analyzing $D_{1}$ via classical number theory.
Proposition 3.1.
The probability that $D_{1}$ does not hold is $o\left(p^{4}\right)$.
@@
Let $\pi(x)$ denote the number of primes less than or equal to $x$. The prime number theorem states that as $x$ tends to $\infty$, $\pi(x)=\frac{x}{\ln x}\left(1+o(1)\right)$. It follows that for any constant $c>1$,
$\pi(cx)-\pi(x)=\frac{(c-1)x}{\ln x}\left(1\pm o(1)\right).$
Let $N(p)$ be the number of primes $q$ such that $f(p)+1\leq q\leq 6f(p)$. Since $f(p)$ tends to infinity as $p$ tends to zero, we have
$N(p)=\frac{\frac{5}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{2}}{\ln\left(\frac{1}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{2}\right)}\left(1\pm o(1)\right)=\frac{\frac{5}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{2}}{\ln\left(\frac{1}{p}\right)\left(1+o(1)\right)}\left(1\pm o(1)\right)\\ =\frac{5}{p}\ln\left(\frac{1}{p}\right)\left(1\pm o(1)\right).$
It follows that
$\mathrm{Pr}\left[\lnot D_{1}\right]=(1-p)^{N(p)}\leq e^{N(p)}=e^{\frac{5}{p}\ln\left(\frac{1}{p}\right)\left(1\pm o(1)\right)}=p^{5\left(1\pm o(1)\right)}=o\left(p^{4}\right).$
Remark.
We could make the bound on $\mathrm{Pr}(\lnot D_{1})$ effective by applying an effective version of the prime number theorem, such as Rosser’s classical result [Ros41] that for all $x\geq 55$,
@@
Next, to analyze the event $D_{2}$, assume that $D_{1}$ holds and define $q$ to be the largest prime between $f(p)$ and $cf(p)$ that is selected as a generator. Let $G$ be the number of selected generators between 1 and $q-1$.
@@
The reason that this may not be obvious is that $q$ itself is a random variable that depends on which generators were selected. Indeed, if we had defined $q$ to be the smallest prime or a random prime rather than the largest prime among the generators between $f(p)+1$ and $cf(p)$, then this statement would not hold.
@@
$\mathrm{E}[G_{q_{i}}|q=q_{j}]=\frac{\mathrm{Pr}\left[(\textup{select }q_{i})\land(q=q_{j})\right]}{\mathrm{Pr}[q=q_{j}]}=\frac{p^{2}(1-p)^{n-j}}{p(1-p)^{n-k}}=p,$
@@
The conditional probability $\mathrm{Pr}[\lnot D_{2}|D_{1}]$ is $O\left(p^{\frac{1}{8}\ln\left(\frac{1}{p}\right)}\right)$. In particular, it decays faster than any fixed power of $p$.
@@
$\mathrm{E}[G]=p(q-1)\geq pf(p)=\left(\ln\left(\frac{1}{p}\right)\right)^{2}.$
On the other hand,
$12\ln q\leq 12\ln\left(\frac{5}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{2}\right)=O\left(\ln\left(\frac{1}{p}\right)\right).$
Thus for sufficiently small $p$, we have $12\ln q<\frac{\mathrm{E}[G]}{2}$. We apply the Chernoff bound
@@
with $\lambda=\frac{\mathrm{E}[G]}{2}$ to obtain
$\mathrm{Pr}[\lnot D_{2}|D_{1}]\leq\mathrm{Pr}\left[G\leq\frac{\mathrm{E}[G]}{2}\right]\leq e^{-\frac{\mathrm{E}[G]}{8}}\leq e^{\frac{-1}{8}\left(\ln\left(\frac{1}{p}\right)\right)^{2}}=p^{\frac{1}{8}\ln\left(\frac{1}{p}\right)}.\qed$
Remark.
To get a significantly faster decay in Proposition 3.3 in practice, we could select an appropriate constant $d$ and modify $f(p)$ by replacing both occurrences of $\frac{1}{p}$ by $\frac{d}{p}$. The constant $\frac{1}{8}$ in the exponent would become $\frac{d}{8}$. The tradeoff is that the bound on the Frobenius number in (4) becomes somewhat weaker. We could also apply Chernoff more tightly for appropriately small values of $p$.
We now assume that events $D_{1}$ and $D_{2}$ both hold, so that we can use Theorem 2.1 in order to analyze $D_{3}$.
Proposition 3.4.
The conditional probability $Pr[\lnot D_{3}|D_{1}\land D_{2}]$ is $o\left(p^{4}\right)$.
Proof.
By applying Theorem 2.1 with $b=6$ along with the lower and upper bounds on the prime $q$, we have
$\mathrm{Pr}[\lnot D_{3}|D_{1}\land D_{2}]\leq\frac{12\log_{2}q+3}{q^{4}}\leq\frac{12\log_{2}\left(\frac{6}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{2}\right)+3}{\left(\frac{1}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{2}\right)^{4}}=O\left(\frac{\ln\left(\frac{1}{p}\right)}{\left(\frac{1}{p}\right)^{4}\left(\ln\left(\frac{1}{p}\right)\right)^{8}}\right)$
which is $o\left(p^{4}\right)$. ∎
@@
The probability $\mathrm{Pr}[\lnot D_{3}]$ that the Apéry set process fails is $o\left(p^{4}\right)$.
@@
We calculate
$\mathrm{Pr}[\lnot D_{3}]$ $=\mathrm{Pr}[\lnot D_{1}]+\mathrm{Pr}[D_{1}\land\lnot D_{2}]+\mathrm{Pr}[D_{1}\land D_{2}\land\lnot D_{3}]$
$\leq\mathrm{Pr}[\lnot D_{1}]+\mathrm{Pr}[\lnot D_{2}|D_{1}]+\mathrm{Pr}[\lnot D_{3}|D_{1}\land D_{2}]$
and all three terms are $o\left(p^{4}\right)$ by Propositions 3.1, 3.3, and 3.4. ∎
@@ 4. Proof of the Main Theorem
To prove Theorem 1.6, we will combine Corollary 3.5 with an alternative bound on the Frobenius number when the event $D_{3}$ does not hold, as follows. It is well known ([RGS09, Proposition 2.13]) that for any coprime numbers $n_{1},n_{2}$, the Frobenius number of the semigroup $\langle n_{1},n_{2}\rangle$ is exactly $(n_{1}-1)(n_{2}-1)-1$. In particular, as soon as two consecutive numbers $12n,2n+$ 1 (the first one even) are selected as generators of $\mathcal{S}$, then
@@
$\mathrm{E}[\operatorname{F}(\mathcal{S})|\operatorname{F}(\mathcal{S})\geq u]=\frac{8}{p^{4}}+\frac{4u}{p^{2}}+u^{2}.$
@@
Let $L$ be the smallest number such that $L>u/2$ and such that both $2L$ and $12L+$ 1 are selected as generators of $\mathcal{S}\in\mathcal{S}(p)$. For any given $p$, this random variable is finite with probability one. Also, $L$ depends only on the selection of generators greater than $u$. On the other hand, the event that $\operatorname{F}(\mathcal{S})\geq u$ depends only on the generators less than or equal to $u$, so $L$ is independent of this event. Thus
$\mathrm{E}[\operatorname{F}(\mathcal{S})|\operatorname{F}(\mathcal{S})\geq u]$ $\leq\mathrm{E}[\operatorname{F}\left(\langle 2L,2L+1\rangle\right)]$
$\leq\mathrm{E}[4L^{2}]$ $=\sum_{n=\frac{u}{2}+1}^{\infty}4n^{2}\left(1-p^{2}\right)^{n-\frac{u}{2}}p^{2}$
@@
$=\frac{8}{p^{4}}+\frac{4u}{p^{2}}+u^{2}.\qed$
Let $F$ and $Z$ be random variables that represent the Frobenius number and the embedding dimension, respectively, of a numerical semigroup $\mathcal{S}\in\mathcal{S}(p)$. As we have seen, we can bound the probability of the undesirable outcome $\lnot D_{3}$ by $o\left(p^{4}\right)$ to cancel out the factor of $O\left(\frac{1}{p^{4}}\right)$ in Lemma 4.1. Choose
$u(p)=K\left(\frac{1}{p}\left(\ln\left(\frac{1}{p}\right)\right)^{3}\right)$
where the constant $K$ is as in (4), so that $\operatorname{F}(\mathcal{S})\leq u(p)$ whenever the event $D_{3}$ holds.
@@
To prove the upper bounds on both the expected Frobenius number and the expected embedding dimension, we will condition on whether or not the Frobenius number is bounded by $u(p)$.
For the expected Frobenius number, we apply Corollary 3.5 along with (4) to obtain